# Sampling the space of solutions of an artificial neural network

### v1

@@ III.1.3 The center is not flat
We then studied the geometry of the center of the spiky low energy manifold in more detail using coupled replica simulations, starting with five GD solutions coupled to their barycenter and increasing the harmonic couplings until all the running points converge to a single vector $\boldsymbol{c}$ (Fig. 4a).
@@
In addition, the similarity distribution between centers and anchor points is larger than that between centers and GD solutions. This ordering suggests that the centers lie deep within the bulk of the solution manifold, with anchor points positioned more peripherally and GD solutions even farther away. This “nested overlap” structure has already been observed in the negative perceptron problem, see Annesi et al. (2024).
@@
From this last computational analysis, we conclude that the center of the spiky low–energy manifold is not barrier–free. Although there are low–energy paths connecting all low–energy regions of the structure, these are simple as linear paths. Moreover, there is a small slope of the energy towards the center.
@@ III.2.1 Similarities and differences with the underparametrized regime
We then compared the properties of the space of weights that we found at the interpolation threshold with that of the overparametrized regime.
We performed various HMC simulations at $\alpha=0.2$ for a wide range of temperatures, all of which never indicate a frozen behavior as, similarly to the intermediate–temperature case at $\alpha=1.8$, the intra e inter overlap distribution coincide, see Fig. 5a. Moreover, the energy profile along the linear paths between the sampled states shows a central free energy barrier, both at $\alpha=1.8$ and at most temperatures at $\alpha=0.2$. In the last case, the central barrier disappears for vanishing temperatures and the energy profile assumes a convex shape (Fig. 5b).
@@
Figure 5: a) Upper panel: intra-state (thin line histograms) and the inter-state overlap distribution (thick line histogram) for two simulations close to the interpolation threshold of GD ($\alpha=1.8$) and $T=1.8\cdot 10^{1}$ (intermediate temperature). The three overlapping distributions have been slightly shifted along the $x$–axis to facilitate the comparison). Lower panel: for the same $(\alpha,T)$, we show the binary cross-entropy loss along the mean geodesic curve for points sampled at equilibrium, parametrized by the variable $\gamma\in[0,1]$. b) The same quantities are presented for simulations in the overparametrized regime ($\alpha=0.2$) and $T=6\cdot 10^{-3},2\cdot 10^{-3},6\cdot 10^{-4}$ (intermediate to low temperatures).
@@ III.2.2 The barrier along linear paths between sampled solutions can be computed analytically
We are interested in computing the average training error and training loss landscape on the geodesic path joining $\boldsymbol{w}^{1}$ and $\boldsymbol{w}^{2}$ at fixed squared norm $Q$. This can be obtained as follows. Given $\boldsymbol{w}^{1}$ and $\boldsymbol{w}^{2}$ we define the interpolating weight $\boldsymbol{w}(\gamma)$ with parameter $\gamma\in[0,1]$ as the vector whose components are the linear interpolation of the components of $\boldsymbol{w}^{1}$ and $\boldsymbol{w}^{2}$
@@
[…]
@@
Using a general loss $\widetilde{\mathcal{L}}$ allows us also to both have access to training loss and training error profiles. Equation (27) can be computed by the replica method in the large $N$ limit. In the following we will focus only on the infinite width limit $K\to\infty$ with the ratio $K/N\to 0$, as done in Baldassi et al. (2019). The full calculation is reported in Appendix B. Here we only mention that the result of the calculation, assuming no replica symmetry breaking of the solution space (i.e. in the so called Replica Symmetric ansatz), will depend on simple geometrical quantities. Those are the typical overlap between two weights $\boldsymbol{w}^{a}$, $\boldsymbol{w}^{b}$ extracted from $p_{1}$, $\boldsymbol{w}^{a},\boldsymbol{w}^{b}\sim p_{1}(\boldsymbol{w};\mathcal{D})$ (respectively from $p_{2}$, i.e. $\boldsymbol{w}^{a},\boldsymbol{w}^{b}\sim p_{2}(\boldsymbol{w};\mathcal{D})$), i.e.
@@
as well as the typical overlap $p$ between the endpoints $\boldsymbol{w}^{1}$ and $\boldsymbol{w}^{2}$, which was introduced in equation (26). In the Replica Symmetric ansatz all those overlaps concentrate in the large $N$ limit. Since the weights $\boldsymbol{w}^{1}$ and $\boldsymbol{w}^{2}$ are in general samples of two different probability distributions (22), the overlaps $q_{1}$ and $q_{2}$ can be simply obtained by computing with the replica method the corresponding partition function $Z^{1}_{\mathcal{D}}$ and $Z^{2}_{\mathcal{D}}$. We refer to Baldassi et al. (2019, 2023) for such calculation, but we report for convenience of the reader in Appendix A its outcome. The overlap $p$ is instead slightly more difficult to compute, because it amounts to study an elastically coupled system of weights $\boldsymbol{w}^{1}$, $\boldsymbol{w}^{2}$ in the limit where the coupling is sent to zero, see Annesi et al. (2023) for details. We present in Appendix D an alternative and complementary calculation based on the Franz-Parisi potential Franz and Parisi (1995).
@@ III.2.3 Comparison between theory and simulations
Figure 6: Loss landscape along the (fixed norm) geodesic path interpolating two weight configurations sampled from the Boltzmann distribution (7) with the cross entropy loss, temperature $T=0.02$ and regularization parameter $\beta\lambda=0.02$. The ReLU function is used as the activation function. We have fixed the constrained density to $\alpha=0.2$ and $K/N=0.005$, while increasing $N$ (points). Full line refers to the theoretical prediction, which is reported in equation (62) of the appendix.
@@ III.2.4 The star-shaped property of the solution space
In Figure 8 we consider the case in which the endpoints are sampled from the large $\beta$ limit of the Boltzmann distributions corresponding to the loss (6), but with two different margin $\kappa_{1}$ and $\kappa_{2}$. We consider the first endpoint to be a typical solution of the classification task, i.e. to have margin $\kappa_{1}=0$. Similarly as before, we plot the training error, i.e. ~$\widetilde{\mathcal{L}}(\widetilde{\boldsymbol{w}};\mathcal{D})=\sum_{\mu=1}^{P}\Theta(-y^{\mu}\Delta^{\mu}(\boldsymbol{w}))$, on the geodesic path joining $\boldsymbol{w}^{1}$ to $\boldsymbol{w}_{2}$ for several values of $\kappa_{2}$ and for the same $\alpha$ considered in Figure 7. It can be noticed that the maximum of the barrier is always closer to the less robust solution, and the whole curve monotonically decreases if one keeps increasing $\kappa_{2}$. For $\kappa_{2}>\kappa_{\text{krn}}$ the two solutions become eventually linear mode connected. This means that typical solutions despite not being linear mode connected, are connected by a piecewise path, passing through a solution having a rather large margin $\kappa$. It therefore exists a subset of solutions called kernel, that are geodesically connected to any other solution of the learning task [^1] Indeed if $\kappa_{2}>\kappa_{\text{krn}}$, then not only this solution is linear mode connected to a typical solution with $\kappa_{1}=0$, but also to any other solution with margin $\kappa_{1}>0$ extracted from the Boltzmann measure induced by (6).. This implies that the space of solutions is star-shaped in the overparameterized regime. Similar plots hold for other activation functions; we report the case of the Erf activation function in the appendix. This conclusion is consistent with what was found in reference Annesi et al. (2023) for a non-convex but simpler linear model called the negative perceptron. Recently in Lin et al. (2024); Sonthalia et al. (2024), numerical evidence has been presented suggesting that the solutions space of deep networks possess a star-shaped geometry. We first provide theoretical support for this claim in the case of simple, overparameterized one-hidden-layer networks with general activation functions.
@@
Figure 7: Training error along the geodesic path connecting 2 solutions sampled from the loss function $\ell_{r}(x)\equiv\Theta(\kappa_{r}-x)$, $r=1,2$ with equal margin $\kappa_{1}=\kappa_{2}\equiv\kappa$, $\alpha=0.1$ and with fixed norm $Q=1$. The two panels refers to the choice of the activation function: ReLU activation (left) and sign activation (right panel). The barrier between typical solutions i.e. $\kappa=0$ is strictly non-vanishing; increasing the margin on the sampled solutions the barrier decreases and eventually vanishes for large enough $\kappa$, similarly to the finding of Annesi et al. (2023).
Figure 8: Case where $\ell_{r}(x)\equiv\Theta(\kappa_{r}-x)$, $r=1,2$ and ~$\widetilde{\ell}(x)=\Theta(-x)$, with $\alpha=0.1$ and fixed norm $Q=1$. We here fixed the first endpoint (that is located at $\gamma=1$) to have a margin $\kappa_{1}=0$. The second endpoint ($\gamma=0$) has a variable value of the margin $\kappa_{2}$ (from smaller to larger margin, curves from top to bottom). The left panel corresponds to the ReLU activation, the right one to the sign function. Increasing the robustness $\kappa_{2}$ the training error barrier on the geodesic monotonically decreases. For large enough $\kappa_{2}$ the solutions become geodesically connected, as observed in Annesi et al. (2023) for the negative perceptron model. The solution space is therefore star-shaped, namely there exists a set o solutions (those ones with a very large margin), that are geodesically connected to all other solutions with lower margin.
@@ III.3 The geometrical structure can be robust with respect to highly correlated data
Similarly to the case of random inputs, the energy of the sampled state decreases as we move from the GD solutions to the center (upper panel in Fig. 9b). An important difference is that now the distribution of similarities between the center and the GD solutions partially overlaps with the similarity between the center and the points $\boldsymbol{s}^{\star}$ found by the double ratchet to connect the GD solutions. In turn, this distribution partially overlaps with that of similarity between the different central points (lower panel in Fig. Fig. 9b). This suggests that the spiky manifold has a larger center for correlated data.
Moreover, the energy profile seems to display higher peaks and to be more rugged than the random label case, as apparent from the shape of the energy along the paths generated by the double ratchet (Fig. 9c) and from the irregular relation between the height of the peaks found along the linear paths (compare the upper panel of Fig. 9b with that of Fig. 4b).
@@ IV Discussion and Conclusions
The symmetric, hollow structure of states at intermediate temperature is similar to that displayed by the system in the overparametrized regime, suggesting that the two manifolds are quite similar. The main difference between the overparametrized regime and that at the interpolation threshold is at low temperature; while in the latter case we have the spiky shape of the manifold discussed above, in the former they are more spherical, located at the center of the manifold.
In the overparametrized regime we have also resorted to replica computation to study both the loss and the training error landscape on the linear interpolation between two weights sampled with different Boltzmann probability distributions. Our work shows that the training error manifold is star-shaped: it exists a subset of robust, solutions having a large margin $\kappa$ that are linear mode connected to solutions having lower (even zero) margin, similarly to what was found in Annesi et al. (2023). Typical low temperature weights extracted from the Boltzmann measure equipped with the cross entropy loss function tend to focus on the inner core of the star-shaped manifold, that we called geodesically convex component, as any two solutions in this region are linear mode connected. This result also agrees with numerical simulations. Our theory also reproduces the non-monotonic behavior of the energy along the linear interpolation between Boltzmann samples at larger temperatures.
@@ A.1 Large $\beta$ limit
$-f$ $\equiv\lim_{\beta\to\infty}\phi=\mathrm{extr}_{\delta q,Q}\left[\mathcal{G}_{S}(\delta q,Q)+\alpha\mathcal{G}_{E}(\delta q,Q)\right]$ (37a)
@@
$\begin{split}z_{\star}(z_{0})&=\operatorname*{argmax}_{z_{1}}\left[-\frac{z_{1}^{2}}{2}-\ell_{1}\left(\sqrt{\Delta_{Q}(Q)-\Delta_{Q}(0)}\,z_{0}+\sqrt{\Delta_{Q}^{\prime}(Q)\delta q_{1}}\,z_{1}\right)\right]\end{split}$ (38)
@@ Appendix B Loss Landscape on the linear interpolation between weights
we have here focused on a particular pattern $\mu$, since in the thermodynamic limit all input patterns will give the same contribution on average.

### v2

@@ III.1.3 The center is not flat
We then studied the geometry of the center of the spiky low energy manifold in more detail using coupled replica simulations, starting with five GD solutions coupled to their barycenter and increasing the coupling term $\gamma$ (Eq. (21)) until all the running points converge to a single vector $\boldsymbol{c}$ (Fig. 4a).
@@
In addition, the similarity distribution between centers and anchor points is larger than that between centers and GD solutions. This ordering suggests that the centers lie deep within the bulk of the solution manifold, with anchor points positioned more peripherally and GD solutions even farther away. This “nested overlap” structure has already been observed in the negative perceptron problem, see ref. Annesi et al. (2025).
@@
From this last computational analysis, we conclude that the center of the spiky low–energy manifold is not barrier–free. Although there are low–energy paths connecting all low–energy regions of the structure, these are not as simple as linear paths. Moreover, there is a small slope of the energy towards the center.
@@ III.2.1 Similarities and differences with the underparametrized regime
We then compared the properties of the space of weights that we found at the interpolation threshold with that of the overparametrized regime $\alpha=0.2$ (i.e. $P=200$). The presented results are obtained from different simulations in the stationary regime and averaged over several dataset instances.
We performed various HMC simulations for a wide range of temperatures, all of which never indicate a frozen behavior, as, similarly to the intermediate–temperature case at $\alpha=1.8$, the intra e inter overlap distribution coincide, see Fig. 5a. Moreover, the energy profile along the linear paths between the sampled states shows a central free energy barrier, both at $\alpha=1.8$ and at most temperatures at $\alpha=0.2$. In the last case, the central barrier disappears for vanishing temperatures and the energy profile assumes a convex shape (Fig. 5b).
@@
Figure 5: a) The thin–line histogram in the upper panel shows two intra-state overlap distributions $\rho(q)$ associated with two intermediate-temperature ($T=1.8\cdot 10^{1}$) simulations starting from different GD solutions close to the interpolation threshold ($\alpha=1.8$). The thick–line histogram stands for the corresponding inter-state distribution computed between the two sets of samples states. The three overlapping distributions have been slightly shifted along the $x$–axis to facilitate the comparison. Lower panel: for the same $(\alpha,T)$, we show the binary cross-entropy loss along the mean geodesic curve for points sampled at equilibrium, parametrized by the variable $\gamma\in[0,1]$. b) The same quantities are presented for simulations in the overparametrized regime ($\alpha=0.2$) at $T=6\cdot 10^{-3},2\cdot 10^{-3},6\cdot 10^{-4}$ (intermediate to low temperatures).
@@ III.2.2 The barrier along linear paths between sampled solutions can be computed analytically
We are interested in computing the average training error (remind Eq. (3) and the discussion in section II.1) and training loss landscape on the geodesic path joining $\boldsymbol{w}^{1}$ and $\boldsymbol{w}^{2}$ at fixed squared norm $Q$. This can be obtained as follows. Given $\boldsymbol{w}^{1}$ and $\boldsymbol{w}^{2}$ we define the interpolating weight $\boldsymbol{w}(\gamma)$ with parameter $\gamma\in[0,1]$ as the vector whose components are the linear interpolation of the components of $\boldsymbol{w}^{1}$ and $\boldsymbol{w}^{2}$
@@
$p\equiv\frac{1}{N}\sum_{li}w_{li}^{1}w_{li}^{2}$ (26)
which corresponds to the overlap between $\boldsymbol{w}^{1}$ and $\boldsymbol{w}^{2}$.
@@
[…]
Equation (27) can be computed by the replica method in the large $N$ limit. In the following we will focus only on the infinite width limit $K\to\infty$ with the ratio $K/N\to 0$, as done in Baldassi et al. (2019). The full calculation is reported in Appendix B. Here we only mention that the result of the calculation, assuming no replica symmetry breaking of the solution space (i.e. in the so called Replica Symmetric ansatz), will depend on simple geometrical quantities. Those are the typical overlap between two weights $\boldsymbol{w}^{a}$, $\boldsymbol{w}^{b}$ extracted from $p_{1}$, $\boldsymbol{w}^{a},\boldsymbol{w}^{b}\sim p_{1}(\boldsymbol{w};\mathcal{D})$ (respectively from $p_{2}$, $\boldsymbol{w}^{a},\boldsymbol{w}^{b}\sim p_{2}(\boldsymbol{w};\mathcal{D})$), i.e.
@@
as well as the typical overlap $p$ between the endpoints $\boldsymbol{w}^{1}$ and $\boldsymbol{w}^{2}$, which was introduced in equation (26). Due to the concentration of the Gibbs measure in the large $N$ limit $p$ can be written as
Since the weights $\boldsymbol{w}^{1}$ and $\boldsymbol{w}^{2}$ are in general samples of two different probability distributions (22), the overlaps $q_{1}$ and $q_{2}$ can be simply obtained by computing with the replica method the corresponding partition function $Z^{1}_{\mathcal{D}}$ and $Z^{2}_{\mathcal{D}}$. We refer to Baldassi et al. (2019, 2023) for such calculation, but we report for convenience of the reader in Appendix A its outcome. The overlap $p$ is instead slightly more difficult to compute, because it amounts to study an elastically coupled system of weights $\boldsymbol{w}^{1}$, $\boldsymbol{w}^{2}$ in the limit where the coupling is sent to zero, see Annesi et al. (2023) for details. We present in Appendix D an alternative and complementary calculation based on the Franz-Parisi potential Franz and Parisi (1995).
@@ III.2.3 Comparison between theory and simulations
Figure 6: Loss landscape along the (fixed norm) geodesic path interpolating two weight configurations sampled from the Boltzmann distribution (7) with the cross entropy loss, temperature $T=0.02$ and regularization parameter $\beta\lambda=0.02$. The ReLU function is used as the activation function. We have fixed the constrained density to $\alpha=0.2$ and $K/N=0.005$, while increasing $N$ (points). Full line refers to the theoretical prediction, which is reported in Eq. (64) of the appendix.
@@ III.2.4 The star-shaped property of the solution space
In Figure 8 we consider the case in which the endpoints are sampled from the large $\beta$ limit of the Boltzmann distributions corresponding to the loss (6), but with two different margin $\kappa_{1}$ and $\kappa_{2}$. We consider the first endpoint to be a typical solution of the classification task, i.e. to have margin $\kappa_{1}=0$. Similarly as before, we plot the training error, i.e. ~$\widetilde{\mathcal{L}}(\widetilde{\boldsymbol{w}};\mathcal{D})=\sum_{\mu=1}^{P}\Theta(-y^{\mu}\Delta^{\mu}(\boldsymbol{w}))$, on the geodesic path joining $\boldsymbol{w}^{1}$ to $\boldsymbol{w}_{2}$ for several values of $\kappa_{2}$ and for the same $\alpha$ considered in Figure 7. It can be noticed that the maximum of the barrier is always closer to the less robust solution, and the whole curve monotonically decreases if one keeps increasing $\kappa_{2}$. For $\kappa_{2}>\kappa_{\text{krn}}$ the two solutions become eventually linear mode connected. This means that typical solutions despite not being linear mode connected, are connected by a piecewise path, passing through a solution having a rather large margin $\kappa$. It therefore exists a subset of solutions called kernel, that are geodesically connected to any other solution of the learning task [^2] Indeed if $\kappa_{2}>\kappa_{\text{krn}}$, then not only this solution is linear mode connected to a typical solution with $\kappa_{1}=0$, but also to any other solution with margin $\kappa_{1}>0$ extracted from the Boltzmann measure induced by (6).. This implies that the space of solutions is star-shaped in the overparameterized regime. Similar plots hold for other activation functions; we report the case of the Erf activation function in the appendix. This conclusion is consistent with what was found in reference Annesi et al. (2023) for a non-convex but simpler linear model called the negative perceptron. Recently in Lin et al. (2024); Sonthalia et al. (2024), numerical evidence has been presented suggesting that the solutions space of deep networks possess a star-shaped geometry. We first provide theoretical support for this claim in the case of simple, overparameterized one-hidden-layer networks with general activation functions. Analyzing how this structure evolves in the underparameterized regime would necessitate accounting for Replica Symmetry Breaking effects. This study is therefore left to future work.
@@
Figure 7: Training error along the geodesic path connecting 2 solutions sampled from the loss function $\ell_{r}(x)\equiv\Theta(\kappa_{r}-x)$, $r=1,2$ with equal margin $\kappa_{1}=\kappa_{2}\equiv\kappa$, $\alpha=0.1$ and with fixed norm $Q=1$. The two panels refers to the choice of the activation function: ReLU activation (left) and sign activation (right panel). The barrier between typical solutions i.e. $\kappa=0$ is strictly non-vanishing; increasing the margin on the sampled solutions the barrier decreases and eventually vanishes for large enough $\kappa$, similarly to the finding of ref.  Annesi et al. (2023).
Figure 8: Case where $\ell_{r}(x)\equiv\Theta(\kappa_{r}-x)$, $r=1,2$ and ~$\widetilde{\ell}(x)=\Theta(-x)$, with $\alpha=0.1$ and fixed norm $Q=1$. We here fixed the first endpoint (that is located at $\gamma=1$) to have a margin $\kappa_{1}=0$. The second endpoint ($\gamma=0$) has a variable value of the margin $\kappa_{2}$ (from smaller to larger margin, curves from top to bottom). The left panel corresponds to the ReLU activation, the right one to the sign function. Increasing the robustness $\kappa_{2}$ the training error barrier on the geodesic monotonically decreases. For large enough $\kappa_{2}$ the solutions become geodesically connected, as observed in ref.  Annesi et al. (2023) for the negative perceptron model. The solution space is therefore star-shaped, namely there exists a set o solutions (those ones with a very large margin), that are geodesically connected to all other solutions with lower margin.
@@ III.3 The geometrical structure can be robust with respect to highly correlated data
Similarly to the case of random inputs, the energy of the sampled state decreases as we move from the GD solutions to the center (upper panel in Fig. 9b). An apparent difference between the two cases is the increase in the overlap between the similarity distributions computed among $c$ states and among the anchor points $\boldsymbol{s}^{\star}$ found by the double ratchet to connect the GD solutions (lower panel in Fig. Fig. 9b). Also, the distribution of similarities between center vectors spreads over lower values as compared to the random–input case (lower panel in Fig. Fig. 9b and lower panel in Fig. Fig. 4b, respectively). This suggests that the center of the spiky manifold is wider and closer to the out-of-equilibrium GD solutions for correlated data.
Moreover, the energy profile seems to display higher peaks and to be more rugged than the random label case, as apparent from the shape of the energy along the paths generated by double ratchet simulations connecting $\boldsymbol{s}$ and $\boldsymbol{c}$ vectors (Fig. 9c, blue lines) and from the irregular relation between the height of the peaks found along the linear paths and the position in the solutions manifold (compare the upper panel of Fig. 9b with that of Fig. 4b).
@@ IV Discussion and Conclusions
The symmetric, hollow structure of states at intermediate temperature is similar to that displayed by the system in the overparametrized regime, suggesting that the two manifolds are quite similar. The main difference between the overparametrized regime and that at the interpolation threshold is at low temperature; while in the latter case we have the spiky shape of the manifold discussed above, in the former the solutions are more spherically distributed, and located at the center of the manifold.
In the overparametrized regime we have also resorted to replica computation to study both the loss and the training error landscape on the linear interpolation between two weights sampled with different Boltzmann probability distributions. Our work shows that the training error manifold is star-shaped: it exists a subset of robust, solutions having a large margin $\kappa$ that are linear mode connected to solutions having lower (even zero) margin, similarly to what was found in ref.  Annesi et al. (2023). Typical low temperature weights extracted from the Boltzmann measure equipped with the cross entropy loss function tend to focus on the inner core of the star-shaped manifold, that we called geodesically convex component, as any two solutions in this region are linear mode connected. This result also agrees with numerical simulations. Our theory also reproduces the non-monotonic behavior of the energy along the linear interpolation between Boltzmann samples at larger temperatures.
@@
Code availability
The code used for the numerical calculations is freely available on Github.
@@ A.1 Large $\beta$ limit
$-f$ $\equiv\lim_{\beta\to\infty}\frac{\phi}{\beta}=\mathrm{extr}_{\delta q,Q}\left[\mathcal{G}_{S}(\delta q,Q)+\alpha\mathcal{G}_{E}(\delta q,Q)\right]$ (38a)
@@
$\begin{split}z_{\star}(z_{0})&=\operatorname*{argmax}_{z_{1}}\left[-\frac{z_{1}^{2}}{2}-\ell\left(\sqrt{\Delta_{Q}(Q)-\Delta_{Q}(0)}\,z_{0}+\sqrt{\Delta_{Q}^{\prime}(Q)\delta q}\,z_{1}\right)\right]\end{split}$ (39)
@@ Appendix B Loss Landscape on the linear interpolation between weights
we have here focused on a particular pattern $\mu_{\star}$, since in the thermodynamic limit all input patterns will give the same contribution on average.